give the logarithms of the following.
1. log 1000
2. log 3.20
3. log(1000)^1/2 <--- thats the whole exponent the 1/2

Answer by Alan3354(31538)

give the logarithms of the following.
1. log 1000
1000 = 10^3 --> log(1000) = 3
----------------
2. log 3.20 =~ 0.50515
------------------
3. log(1000)^1/2
= 1/2 log(1000) = 3/2 or 1.5